# Quantum querying based on multicontrolled Toffoli gates for causal Feynman loop configurations and directed acyclic graphs

### v1

@@ 4 Efficiency assessment
In this section we assess the efficiency of the current quantum algorithm by comparing the implementation results with the quantum algorithm and the multiloop topologies analysed in Ref. Ramirez-Uribe:2021ubp . These multiloop topologies are depicted in Fig. 1 considering one edge in each set. The results are summarised in Table 2 which displays the number of qubits required for the implementation, the theoretical quantum circuit depth, the related mixing angle and the number of causal configurations.
@@
Fig. eloops (edges) Qubits $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$ $\begin{subarray}{c}\text{Causal}\\ \text{states}\end{subarray}$
1A one ($3$) $\phantom{0}5~{}|~{}\phantom{0}7$ $\phantom{0}6~{}|~{}18$ $37.7^{\circ}$ $0.84$ $3$
1B two ($5$) $\phantom{0}9~{}|~{}14$ $12~{}|~{}21$ $32.0^{\circ}$ $0.99$ $9$
1C three ($6$) $11~{}|~{}19$ $18~{}|~{}24$ $25.7^{\circ}$ $0.95$ $12$
1D four(c) ($8$) $13~{}|~{}25$ $16~{}|~{}23$ $33.5^{\circ}$ $0.97$ $39$
1E, 1F four(t,s) ($9$) $15~{}|~{}28$ $20~{}|~{}25$ $26.5^{\circ}$ $0.97$ $102$
1G four(u) ($9$) $19~{}|~{}33$ $32~{}|~{}28$ $28.3^{\circ}$ $0.99$ $115$
@@
Table 2: Resources required and quantum circuit depth of the quantum algorithms used to analyse the multiloop topologies in Fig. 1. The first number in the second and third columns is from the quantum algorithm proposed in this work, whereas the second number corresponds to Ref. Ramirez-Uribe:2021ubp . The columns with a single entry have the same value for both algorithms. The number in the last column is the number of causal configurations.
@@
The columns of total number of qubits and quantum circuit depth have two numbers: the first one gives the data on the current approach and, the second one the information of the algorithm using binary clauses. The remaining columns are: the initial mixing angle, an indicator for the appropriate selection of the iteration number and the number of causal states found by the algorithm. From Table 2, we observe that in all cases the initial mixing angle is close to $\pi/6$, suggesting the application of one iteration only; from the fifth column we have that the value of the mixing angle after the first iteration approaches to $\pi/2$, the desired probability amplification; with $\sin^{2}(\theta_{t})\sim 1$.
@@
Moving on to the number of qubits, the quantum algorithm proposed in this work lowers significantly the total number of qubits required for the implementation, going from a two qubits reduction in the case of one eloop with three vertices, to a remarkable fourteen qubits reduction in the $u$-channel at four eloops. Regarding the theoretical quantum circuit depth, we notice that the new approach gives smaller values for every topology, with the only exception of the $u$-channel, which is a nonplanar diagram. Delving in the $u$-channel case, it is important to take into account that the values shown in Table 2 corresponds to a configuration with one edge per set. The scenario with additional edges per set does not change the quantum circuit depth of the proposed algorithm whereas the quantum circuit depth of the algorithm using binary clauses increases as additional edges per set are included. Therefore, the behaviour of the quantum circuit depth comparative between both algorithms changes when we start considering more than one-edge-per-set complex multiloop diagrams, giving to the new algorithm an important advantage.
a) Toffoli gate
b) Transpiled Toffoli gate.
Figure 4: a) Quantum circuit representation for the Toffoli gate. b) Equivalent quantum circuit of the Toffoli gate transpiled to the ibmq_belem device from IBMQ.
@@
Continuing with the analysis of the quantum circuit depth, we observe in Table 2 a non-intuitive result, the quantum circuit depth for the three-eloop diagram is greater than the quantum circuit depth for the four-eloop diagram with one four-particle interaction vertex (also called N3MLT in Ref. Ramirez-Uribe:2020hes ). To better understand this behaviour we compare the eloop clauses needed for the implementation of both topologies to contrast the presence of multicontrolled Toffoli gates acting on noncommon qubits along the circuits. The eloop clauses required for the three-eloop topology are
@@
For the three-eloop topology, the eloop clauses in Eq. (4) are translated to the quantum circuit into six multicontrolled Toffoli gates and their required CNOT gates. The clauses $a_{0}^{(3)}$ and $a_{1}^{(3)}$ validate only one direction of the subeloop, requiring a single multicontrolled Toffoli clause, while $a_{2}^{(3)}$ and $a_{3}^{(3)}$ test the subeloop in both directions, needing two multicontrolled Toffoli clauses each. In this case, we have all the Toffoli gates acting in at least one common edge, therefore, they have to be applied in different time-steps.
Regarding the four-eloop case, the eloop clauses in Eq. 4 are implemented into the quantum circuit with eight multicontrolled Toffoli gates together with the corresponding CNOT gates; $a_{0}^{(4)}$ and $a_{2}^{(4)}$ need a single multicontrolled Toffoli clause, whereas $a_{1}^{(4)}$, $a_{3}^{(4)}$ and $a_{4}^{(4)}$ require two multicontrolled Toffoli clauses each. The edge configuration of this topology generates several multicontrolled Toffoli gates acting on noncommon qubits allowing to perform some of them at the same time-step.
An important feature of multicontrolled Toffoli gates to be discussed is transpilation. The transpilation process consists of compiling a given quantum circuit to match the specific topology and native gate set of a particular quantum device hardware, as well as its optimization in order to run on the noisy intermediate scale quantum (NISQ) era computers. In the quantum circuit model, the high level quantum circuits are built with unitary operators that are implemented as quantum gates. However, when running them on a real quantum hardware, they must undergo the decomposition of the unitary quantum gates into the native gate set of the target quantum device, and optimize them to reduce the noise effects on the results output.
We illustrate the impact on the quantum circuit depth for the transpilation of a Toffoli gate. The target device is ibmq_belem which has a set of native gates consisting of $I$, $R_{Z}$, $SX$, $X$, CNOT and a reset gate. The Toffoli gate in Fig. (4a) has a theoretical quantum depth of 1 while its transpilation process, Fig. (4b) results in a larger quantum depth of 19. We can see that for each Toffoli gate in the quantum circuit transpilation is costly increasing the quantum depth considerably. Therefore, this has an appreciable impact on the quantum noise and gate errors.
@@
Figure 5: Quantum circuit for the four-eloop topology with eight edges (Fig. 1D).
For example, the four-eloop topology with eight edges is implemented with the quantum circuit of Fig. (5) and requires fourteen qubits, while the binary-clause approach requires twenty five qubits. The transpilation process for the multicontrolled-Toffoli approach requires about 900 native gates on the ibmq_melbourne quantum device while with the binary-clause approach grows to around 1600 native gates in the ibmq_manila quantum device.
@@
5 Multiloop topologies with multiple number of edges in each set
eloops $|{q}\rangle$ $|{a}\rangle$ Total Qubits
@@ 5 Multiloop topologies with multiple number of edges in each set
Table 3: Number of qubits required for an arbitrary three- and four-eloop configuration and the five-eloop configuration in Fig. 6E. The total number of qubits includes the qubit storing the oracle’s marker.
@@
In the previous section, we have shown a significant reduction on the total number of qubits required for the implementation of the quantum algorithm proposed in this work. This improvement allows to explore more complex topologies than those affordable with the algorithm presented in Ref. Ramirez-Uribe:2021ubp . In this section, we inquire into the implementation of the proposed algorithm to three-, four- and five-eloop topologies; the number of qubits required for an arbitrary number of edges per set is shown in Table 3. The particular cases presented are shown in Fig. 6; the number of eloop clauses to test in the oracle operator, the total number of qubits needed for the implementation, the quantum circuit depth and the mixing angle are summarised in Table 4.
@@ 5.1 Three eloops with nine and twelve edges
Figure 7: Quantum circuit for the three-eloop topology with twelve edges (Fig. 6A).
At three eloops, we work with the selected multiloop topologies shown in Figs. 6A and 6B. The first topolology (Fig. 6A) considers two edges in the external sets and one edge in the internal sets, and the Boolean functions are given by,
@@
It requires the four eloop clauses in Eq. (4).
For the second topology (Fig. 6B), we take two edges in all the sets, and the Boolean functions are given by,
@@
It requires in addition to the eloop clauses in Eq. (4), the following eloop clauses to test acyclicity over four sets of edges
@@
The total number of qubits required for the algorithm implementation of the configurations in Eq. (12) and Eq. (13) is fourteen and twenty one respectively, and the quantum circuit depth results in eighteen and thirty two, respectively. In comparison with the binary-clause approach, these configurations require twenty five and thirty six qubits, respectively; the latter exceeding the number of qubits available of current simulators. The quantum circuit depth for three eloops with nine edges is thirty two. The quantum circuit implemented for the configuration in Eq. (13) is shown in Fig. 7.
Regarding the eloop clauses, we have observed that the topology considering multiple edges located exclusively in the external sets do not enlarge the minimum number of clauses to test, whereas including two edges in every set requires to consider the three subloops involving four sets of edges.
@@ 5.2 Four eloops with twelve and sixteen edges
Figure 8: Quantum circuit for the four-eloop topology in Fig. 6C. The output is equal to 1199 causal configurations out of 4096 states.
The four-eloop complexity examples to analyze are those depicted in Figs. 6C and 6D. The first topology (Fig. 6C) considers two edges in the external sets and one edge in the internal sets, and are given by
@@
The second topology (Fig. 6D), consists in two edges in every set,
@@
The complete set of clauses for the four-eloop topology of Fig. 6D, together with the clauses from Eq. (4) are given by
@@
The implementation of the algorithm for these cases requires a total number of eighteen and thirty qubits respectively; the quantum circuit depth is sixteen and fourty six. The quantum circuit implemented for the topology configuration in Eq. (15) is shown in Fig. 8.
The topology associated to Eq. (16) requires thirty qubits which is within the capacities of Qiskit, nevertheless the number of shots needed to select all the causal configurations is beyond the Qiskit simulator capacities. Similar to the previous cases of three eloops, the topology associated to Eq. (15) do not need to test additional eloop clauses besides the minimum requirement. For the topology configuration in Eq. (16) with two edges in all the sets, it is necessary to check clauses with two and three subloops or equivalently, with four and five sets of edges, giving a total of thirteen eloop clauses for all the causal configurations.
Fig. eloops(edges) $|{e}\rangle$ $|{a}\rangle$ $\begin{subarray}{c}\text{Total}\\ \text{Qubits}\end{subarray}$ $\begin{subarray}{c}\text{Toffoli}\\ \text{Gates}\end{subarray}$ $\begin{subarray}{c}\text{NOT}\\ \text{Gates}\end{subarray}$ $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$ $\begin{subarray}{c}\text{Causal}\\ \text{States}\end{subarray}$
6A three ($9$) $9$ $4$ $14$ $14$ $45$ $18$ $35.2^{\circ}$ $0.93$ $170$
6B three($12$) $13$ $7$ $21$ $21$ $48$ $32$ $28.0^{\circ}$ $0.99$ $1804$
6C four(c)($12$) $12$ $5$ $18$ $37$ $66$ $16$ $32.8^{\circ}$ $0.98$ $1199$
6D four(c)($16$) $17$ $13$ $30$ $37$ $85$ $46$ $27.7^{\circ}$ $0.98$ $28343$
6E five(c)($10$) $10$ $6$ $17$ $25$ $37$ $42$ $28.9^{\circ}$ $0.99$ $240$
Table 4: Number of qubits required in the registers $|{e}\rangle$ and $|{a}\rangle$, respectively, for the algorithm implementation of the multiloop topologies depicted in Fig. 6, number of gates needed, quantum circuit depth, the mixing angle, and the number of causal states obtained from the algorithm.

### v2

@@ 4 Efficiency assessment
In this section we assess the efficiency of the proposed quantum algorithm by comparing the implementation results with the quantum algorithm and the multiloop topologies analysed in Ref. [34]. These multiloop topologies are depicted in Fig. 1 considering one edge in each edge set $s_{i}$. The results are summarised in Table 2 which displays the number of qubits required for the implementation, the theoretical quantum circuit depth, the related mixing angle and the number of causal configurations.
@@
Fig. eloops (edges) Qubits $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$ $\begin{subarray}{c}\text{Causal}\\ \text{states}\end{subarray}$ $\begin{subarray}{c}\text{Total}\\ \text{states}\end{subarray}$
1A one ($3$) $\phantom{0}5~{}|~{}\phantom{0}7$ $\phantom{0}6~{}|~{}18$ $37.7^{\circ}$ $0.84$ $3$ $8$
1B two ($5$) $\phantom{0}9~{}|~{}14$ $12~{}|~{}21$ $32.0^{\circ}$ $0.99$ $9$ $32$
1C three ($6$) $11~{}|~{}19$ $18~{}|~{}24$ $25.7^{\circ}$ $0.95$ $12$ $64$
1D four(c) ($8$) $14~{}|~{}25$ $16~{}|~{}23$ $33.5^{\circ}$ $0.97$ $39$ $256$
1E, 1F four(t,s) ($9$) $15~{}|~{}28$ $20~{}|~{}25$ $26.5^{\circ}$ $0.97$ $102$ $512$
1G four(u) ($9$) $19~{}|~{}33$ $32~{}|~{}28$ $28.3^{\circ}$ $0.99$ $115$ $512$
@@
Table 2: Quantum resources required and quantum circuit depth of the quantum algorithms used to analyse the multiloop topologies in Fig. 1. The first number in the third and fourth columns is from the quantum algorithm proposed in this work, whereas the second number corresponds to Ref. [34]. The columns with a single entry have the same value for both algorithms. The numbers in the last two columns are the number of causal configurations and the total number of states, respectively.
@@
The columns of total number of qubits and quantum circuit depth have two numbers: the first one gives the data on the current approach and, the second one the information of the quantum algorithm using binary clauses. The remaining columns are: the initial mixing angle, an indicator for the appropriate selection of the iteration number and the number of causal states. From Table 2, we observe that in all cases the initial mixing angle is close to $\pi/6$, suggesting the application of only one iteration; from the fifth column we have that the value of the mixing angle after the first iteration approaches $\pi/2$, leading to the optimal probability amplification, $\sin^{2}(\theta_{t})\sim 1$ with $t=1$.
@@
Counting the number of qubits needed, the proposed quantum algorithm lowers significantly the total number of qubits required for the implementation with respect to a binary-clause algorithm, going from a reduction of two qubits in the case of one eloop with three vertices, to a reduction of fourteen qubits in the $u$-channel at four eloops. Regarding the theoretical quantum circuit depth, we notice that the new approach yields smaller values for all multiloop topologies, with the only exception of the four-eloop $u$-channel, which is a nonplanar diagram. It is important to note that the values shown in Table 2 correspond to multiloop configurations with one edge per edge set. For these configurations it is sufficient to satisfy the Boolean clauses in Eq. (6) for sub-eloops involving a small number of edge sets, while longer strings of edge sets need to be checked when multiple edges in the edge sets are considered. However, the number of eloop clauses reaches a maximum that depends on the multiloop topology and does not scale with the number of edges (see Table 4), while the number of binary clauses scales with the number of edges. In Section 5, we study multiloop topologies with multiple edges in each edge set, which require testing the maximum number of eloop clauses.
Figure 5: Quantum circuit for the three-eloop topology with six edges (Fig. 1C). The output is 12 causal configurations out of 64 states.
@@
Continuing with the analysis of the quantum circuit depth, we observe in Table 2 that the quantum circuit depth for the three-eloop diagram (Fig. 1C) is greater than the quantum circuit depth for the four-eloop diagram with one four-particle interaction vertex (Fig. 1D, also called N3MLT in Ref. [68]). To better understand this behaviour, we compare the eloop clauses needed to implement the two topologies, and analyse the presence of multicontrolled Toffoli gates acting on noncommon qubits along the circuits. The eloop clauses required for the three-eloop topology are
@@
[…]
In the analysis of the four-eloop scenario, Fig. 1D, the eloop clauses specified in Eq. 4 are implemented into the quantum circuit with eight multicontrolled Toffoli gates: a single multicontrolled Toffoli gate for $a_{0}^{(4)}$ and $a_{2}^{(4)}$, and two multicontrolled Toffoli gates for $a_{1}^{(4)}$, $a_{3}^{(4)}$, and $a_{4}^{(4)}$. In addition, the oracle operator requires ten NOT gates, as shown in Fig. 6. As for the three-eloop case, these oracle operations are also applied in reversed order. The edge configuration of this topology allows several multicontrolled Toffoli gates to act on distinct qubits. This arrangement enables the simultaneous execution of certain operations within the same time-step, which explains the lower quantum circuit depth.
@@
Figure 6: Quantum circuit for the four-eloop topology with eight edges (Fig. 1D). The output is 39 causal configurations out of 256 states.
An important feature of multicontrolled Toffoli gates to be discussed is transpilation. The transpilation process consists of compiling a given quantum circuit to match the specific topology and native gate set of a particular quantum device hardware, as well as its optimisation in order to run on noisy intermediate scale quantum (NISQ) era computers. In the quantum circuit model, high-level quantum circuits are constructed using unitary operators that are implemented as quantum gates. When deploying these circuits on actual quantum hardware, however, they must undergo a decomposition of the unitary quantum gates into the native gate set of the target quantum device. Additionally, optimisation processes are employed to minimise the noise effects and enhance the accuracy of the output results.
The impact of transpilation for a Toffoli gate on the quantum circuit depth is illustrated in Fig 7. The Toffoli gate in Fig 7 operates on three qubits: two control qubits and one target qubit. For our analysis, we employed the 5-qubit quantum processor ibmq_belem (which is now retired [^3] https://docs.quantum.ibm.com/guides/retired-qpus; however, a simulated fake backend [^4] https://docs.quantum.ibm.com/api/qiskit-ibm-runtime/fake_provider remains accessible). This quantum processor contains a native gate set consisting on $I$, $R_{Z}$, $SX$, $X$, CNOT and reset gates. The theoretical quantum circuit depth of the Toffoli gate shown in Fig. 7a is one, however its transpilation for the quantum hardware results in a quantum circuit depth of nineteen as illustrated in Fig. 7b. The transpilation of each Toffoli gate increases the quantum circuit depth and therefore has an impact on quantum noise and gate errors.
(a) Toffoli gate
(b) Transpiled Toffoli gate
Figure 7: (a) Quantum circuit representation for the Toffoli gate. (b) Equivalent quantum circuit of the Toffoli gate transpiled to the ibmq_belem device from IBMQ.
Four-eloop topology of Fig. 1D
Quantum device (qubits) Transpiled quantum circuit depth Two-qubit gates
BC MCX BC MCX
FakeMelbourneV2 (14) not enough qubits 1909 not enough qubits 1562
FakeMumbaiV2 (27) 1495 2163 1484 1762
ibm_brisbane (127) 4188 5695 1558 1748
ibm_kyoto (127) 4726 5969 1558 1748
ibm_osaka (127) 4672 5043 1558 1748
ibm_sherbrooke (127) 4411 5334 1558 1748
Table 3: Key metrics for the transpilation of the quantum circuit corresponding to the four-eloop topology of Fig. 1D in different quantum devices for the binary-clause (BC) and multicontrolled Toffoli (MCX) algorithms.
To analyse the impact of transpilation, we focus on the four-eloop topology of Fig 1D. We study the transpiled quantum circuit depth for relevant metrics in different quantum backends, both simulated and real quantum hardware, as shown in Table 3. The transpiled quantum circuit depth is quite backend-dependent, and is larger for the multicontrolled Toffoli algorithm than for the binary-clause algorithm. The former also requires more two-qubit gates. However, for quantum backends with a limited number of available qubits, only the multicontrolled Toffoli algorithm is capable of simulating complex graph topologies.
It is important to highlight that transpilation may involve more qubits than formally required in the design of the quantum circuit. To address this concern we introduce a new quantum metric, the quantum circuit area, which provides a better understanding of algorithmic efficiency in both quantum simulators and quantum hardware. The quantum circuit area is defined as the product of the quantum circuit depth and the number of qubits required in the transpilation. This metric provides a more comprehensive evaluation depending in both requirements, the quantum depth of the circuit and its qubit usage, yielding a more complete measure of the computational complexity of the quantum circuit.
Figure 8: Quantum circuit depth (upper plot) and quantum circuit area (bottom plot) for the four-eloop topology of Fig 1D using the binary-clause (BC) and the multicontrolled Toffoli (MCX) algorithms for various optimisation levels on currently available quantum hardware backends.
Continuing the analysis of the four-eloop topology shown in Fig 1D, we present in Fig. 8 the transpiled quantum circuit depth and the transpiled quantum circuit area as a function of the optimisation levels on different quantum hardware for both algorithms. The results show a similar evolution in terms of the optimisation levels. Despite the fact that the quantum circuit depth in Fig. 8 is higher for the multicontrolled Toffoli algorithm, the quantum circuit area is much smaller compared to the binary-clause algorithm. Therefore, the proposed algorithm manifests a better balance between quantum circuit depth and qubit count, allowing for an improvement in quantum resource efficiency.
@@
5 Multiloop topologies with multiple number of edges in each edge set
In the previous section, we have shown a significant reduction in the total number of qubits required to implement the quantum algorithm proposed in this work. This achievement allows us to explore more complex topologies than those affordable with the algorithm presented in Ref. [34], due to the limitations of the current quantum simulators in terms of the number of qubits. In this section, we inquire into the implementation of the proposed algorithm to involved three-, four- and five-eloop topologies; the number of qubits required for an arbitrary number of edges per edge set is shown in Table 4. The particular cases to explore are illustrated in Fig. 9. The Table 5 shows the total number of qubits and quantum circuit depth comparing the algorithm proposed in this work with the one in Ref. [34]. This table also provides the mixing angle and the probability of the causal states. Additionally, Table 6 compiles the details for the algorithm implementation, number of edges, eloop clauses to test in the oracle operator, and Toffoli and NOT gates required for the oracle operator. It is important to remark that from the multiloop topologies depicted in Fig. 9, only the three-eloop topology with nine edges and the simplest five-eloop are formally reachable with the algorithm in Ref. [34] in the Qiskit quantum simulator.
eloops $|{e}\rangle$ $|{a}\rangle$ Total Qubits
@@ 5 Multiloop topologies with multiple number of edges in each edge set
Table 4: Number of qubits required for an arbitrary three-, four- and five-eloop configuration in Fig. 9. The total number of qubits includes the oracle’s marker.
@@
Fig. eloops(edges) $\begin{subarray}{c}\text{Total}\\ \text{Qubits}\end{subarray}$ $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$
9A three ($9$) $14~{}|~{}25$ $18~{}|~{}27$ $35.2^{\circ}$ $0.93$
9B three($12$) $21~{}|~{}36$ $32~{}|~{}33$ $28.0^{\circ}$ $0.99$
9C four(c)($12$) $18~{}|~{}33$ $16~{}|~{}31$ $32.8^{\circ}$ $0.98$
9D four(c)($16$) $31~{}|~{}50$ $46~{}|~{}53$ $27.7^{\circ}$ $0.98$
9E five(c)($10$) $17~{}|~{}31$ $24~{}|~{}25$ $28.9^{\circ}$ $0.99$
Table 5: Quantum resources required and quantum circuit depth for the quantum algorithm implementation of the multiloop topologies in Fig. 9. The first and second numbers in the third and fourth columns correspond to the algorithm proposed in this work and to the algorithm in Ref. [34], respectively. The mixing angle and the probability of the causal states after amplitude amplification are the same for both algorithms.
Fig. eloops(edges) $|{e}\rangle$ $|{a}\rangle$ $\begin{subarray}{c}\text{Toffoli}\\ \text{Gates}\end{subarray}$ $\begin{subarray}{c}\text{NOT}\\ \text{Gates}\end{subarray}$ $\begin{subarray}{c}\text{Causal}\\ \text{States}\end{subarray}$ $\begin{subarray}{c}\text{Total}\\ \text{states}\end{subarray}$
9A three ($9$) $9$ $4$ $14$ $45$ $170$ $512$
9B three($12$) $13$ $7$ $21$ $48$ $1804$ $8192$
9C four(c)($12$) $12$ $5$ $17$ $66$ $1199$ $4096$
9D four(c)($16$) $17$ $13$ $39$ $85$ $28343$ $131072$
9E five(c)($10$) $10$ $6$ $25$ $37$ $240$ $1024$
Table 6: Number of qubits required in the $|{e}\rangle$ and $|{a}\rangle$ registers for the algorithm implementation of the multiloop topologies depicted in Fig. 9, number of Toffoli and $X$ gates, number of causal states and total number of states.
@@ 5.1 Three eloops with nine and twelve edges
Figure 10: Quantum circuit for the three-eloop topology with twelve edges (Fig. 9B). The output is 1804 causal configurations out of 8192 states. The last qubit in the $|{e}\rangle$ registed is an ancillary qubit used to adjust the initial mixing angle to the optimal value.
At three eloops, we work with the selected multiloop topologies shown in Figs. 9A and 9B. The first three-eloop topolology (Fig. 9A) considers two edges in the external sets of edges and one edge in the internal sets of edges, and the Boolean functions are given by,
@@
Regarding the eloop clauses, it requires the four eloop clauses in Eq. (4).
For the second three-eloop topology (Fig. 9B), we take two edges in all the sets of edges, and the Boolean functions are given by,
@@
It requires, in addition to the eloop clauses in Eq. (4), the following eloop clauses to test acyclicity over four sets of edges
@@
The total number of qubits required to implement the configurations in Eq. (14) and Eq. (15) is fourteen and twenty-one respectively, and the quantum circuit depth results in eighteen and thirty-two, respectively. In the binary-clause approach these configurations require twenty-five and thirty-six qubits, and the quantum circuit depths would be twenty-seven and thirty-three, respectively; the latter exceeding the number of qubits available on the Qiskit simulator. The quantum circuit implemented for the configuration in Eq. (15) is shown in Fig. 10.
Regarding the number of eloop clauses, any three-eloop topology with multiple edges located exclusively in the external sets of edges do not enlarge the minimum number of clauses to be tested, whereas including two or more edges in every edge set requires to test the maximum number of eloops clauses. In fact, thirty-two is the maximum quantum circuit depth that a three-eloop topology can have due to the fact that the quantum circuit depth of the proposed quantum algorithm increases with the number of eloop clauses but it does not scale with additional edges.
@@ 5.2 Four eloops with twelve and sixteen edges
Figure 11: Quantum circuit for the four-eloop topology in Fig. 9C. The output is 1199 causal configurations out of 4096 states.
The four-eloop examples to analyse are those depicted in Figs. 9C and 9D. The first loop topology (Fig. 9C) considers two edges in the external sets of edges and one edge in the internal sets of edges, and are given by
@@
The second loop topology (Fig. 9D), consists in two edges in every edge set,
@@
The complete set of eloop clauses for the four-eloop topology of Fig. 9D, together with the eloop clauses from Eq. (4) are given by
@@
The implementation of the algorithm for these two cases requires a total number of eighteen and thirty thirty-one qubits, respectively, compared to thirty-three and fifty for the binary-clause approach. In terms of the quantum circuit depth, it is sixteen for Fig. 9C and forty-six for Fig. 9D, hence a maximum quantum circuit depth of forty-six for any four-eloop topology independently of the number of edges. Similar to the previous cases of three eloops, the four-eloop topology associated to Eq. (17) does not need to test additional eloop clauses besides the minimum requirement. The quantum circuit for Fig. 9C is shown in Fig. 11.
For the four-eloop topology configuration in Fig. 9D with two edges in all the sets of edges, it is necessary to check additional eloop clauses, considering those associated with sub-eloops generated with four and five edge sets, Eq. (19), giving a total of thirteen eloop clauses for all the causal configurations.